# Clarification and Coinduction of Tambara Functors

### v1

@@ 4.1. Morphisms between coinduction and clarification
Every free polynomial Tambara functor on finitely many generators has level $G/e$ given by a finitely generated free polynomial $\mathbb{Z}$-algebra over by [Bru05, Theorem A]. The free polynomial $k$-algebras are obtained by base-changing the free polynomial Tambara functors along the unit map from the Burnside Tambara functor to $k$.
@@
$\prod_{H\subset G}\mathrm{Coind}_{H}^{G}R_{H}\rightarrow\prod_{H\in\Lambda}\mathrm{Coind}_{H}^{G}R_{H}\mathrm{.}$
@@
$\prod_{g\in H\backslash G/H}\mathrm{Coind}_{H\cap{}^{g}H}^{H}\mathrm{Res}_{H\cap{}^{g}H}^{{}^{g}H}{}^{g}R\cong\mathrm{Res}_{H}^{G}\mathrm{Coind}_{H}^{G}R\rightarrow S\mathrm{.}$
@@
We observe that each item is clarified in order. First, $G$ acts trivially on the $G/e$ level of the first three kinds Tambara functors on the list, so they are clarified. Next, The fixed-point Tambara functor associated to a Galois field extension has bottom level a field, hence the only idempotents are $1$ and 00, which are fixed by $G$. That free polynomial algebras are clarified over a clarified base is the content of Lemma 4.4. Next, we compute $\mathrm{H}^{2*}(e;\mathrm{Z})\cong\mathrm{Ext}_{\mathbb{Z}}^{2*}(\mathbb{Z},\mathbb{Z})$ which is just $\mathbb{Z}$ in degree zero, and $\underline{\pi}_{0}(MU_{G})(G/e)\cong\pi_{0}(MU)\cong\mathbb{Z}$. Any Tambara functor $k$ with $k(G/e)=\mathbb{Z}$ is clarified, so we deduce that the last two list items are clarified.
@@
We will later observe in Proposition 5.5 another natural and important source of coinduced Tambara functors.
@@
$R\rightarrow\prod_{L\in\Lambda_{H}}\mathrm{Coind}_{L}^{G}R_{L}\mathrm{.}$
@@
$\mathrm{Coind}_{L}^{G}(R_{L}(L/e))\rightarrow\phi_{H}(d_{L})\mathrm{Coind}_{H}^{G}R_{H}(H/e)\mathrm{.}$
@@ 4.2. Clarification as a localization
The subcategory $G{-}\mathsf{Tamb}_{\Lambda{-}\mathrm{clar}}$ of $\Lambda$-clarified $G$-Tambara functors is a reflective subcategory of the category of $G$-Tambara functors, i.e. the inclusion has a left adjoint.
@@
$R\cong\prod_{H\notin\Lambda}\mathrm{Coind}_{H}^{G}R_{H}\mathrm{.}$
@@
Proposition 4.13 is a precise formulation of the slogan “$\Lambda$-clarification is given by setting the image of coinduction equal to zero”. If $G=C_{p^{n}}$ the only clarification functors available to us are the $\Lambda_{C_{p^{m}}}$-clarifications, and a $C_{p^{n}}$-Tambara functor maps to zero under such a clarification if and only if it is in the image of coinduction from $C_{p^{m}}$-Tambara functors (assuming the $C_{p^{n}}/e$ level is Noetherian).
@@
We give an application to étale theory. Recall that a morphism of $G$-Tambara functors $k\rightarrow R$ is étale if $R$ is a finitely presented $k$-algebra, $R$ is flat as a $k$-module, and $k\rightarrow R$ is formally smooth [Hil17].
@@
This follows from Theorem 3.11 and the fact that a product of $k$-algebras is étale if and only if each factor is (cf. [Wis25]). For the reader’s convenience, we sketch the argument. Write $R\cong\widehat{R}\times R^{\mathrm{clar}}$. If $R$ is finitely presented and flat over $k$, then it is straightforward to show that the same holds for $R^{\mathrm{clar}}$. It remains to check formal smoothness. One may show that the (Tambara) genuine Kähler differentials of $R$ over $k$ contain the genuine Kähler differentials of $R^{\mathrm{clar}}$ as a summand, hence the vanishing of the former implies vanishing of the latter. ∎
@@ 4.3. Clarified colimits
$\Lambda$-clarification preserves filtered colimits. In other words, filtered colimits in $G{-}\mathsf{Tamb}_{\Lambda{-}\mathrm{clar}}$ are computed levelwise.
@@
Recall that filtered colimits of Tambara functors are computed levelwise. Since colimits in the category of $\Lambda$-clarified objects are computed by applying $\Lambda$-clarification to the corresponding colimit in $G{-}\mathsf{Tamb}$, it suffices to show that any colimit in $G{-}\mathsf{Tamb}$ of a diagram whose objects are all $\Lambda$-clarified $G$-rings is $\Lambda$-clarified.
Let $F:I\rightarrow G{-}\mathsf{Ring}$ be any filtered diagram and let $R$ be the colimit. Suppose $d\in R$ is an idempotent of type $H$ for some $H\subset G$. Then, passing to a large enough index $i\in I$, $d$ is in the image of some $\delta\in F(i)$ under canonical map $F(i)\rightarrow R$. We see that, modulo the kernel of $F(i)\rightarrow R$, $\delta$ satisfies the requisite relations to be a type $H$ idempotent.
Observe that being a type $H$ idempotent amounts to satisfying a finite list of relations: $hd-d=0$ for all $h\in H$, $d\cdot gd=0$ for all $gH\in G/H$, and $d^{2}-d=0$. Thus there exists $j\in I$ and $i\rightarrow j$ such that the induced map $F(i)\rightarrow F(j)$ carries each relation $h\delta-\delta$, $\delta\cdot g\delta$, and $\delta^{2}-\delta$ to zero in $F(j)$ (because these elements map to zero in the colimit $R$). The image of $\delta$ in $F(j)$ is therefore a type $H$ idempotent.
@@
Note that $\sqrt{-2}\otimes 1+1\otimes\sqrt{-2}\mapsto(2\sqrt{-2},0)$. If we give both factors in $\mathbb{Q}(\sqrt{-2})\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-2})\cong\mathbb{Q}(\sqrt{-2})^{\times|C_{2}|}$ the trivial action, this element is fixed, and if we give both factors the Galois action, then the generator for $C_{2}$ acts by $-1$. However, if we give one factor the trivial action and the other factor the Galois action, then we observe that the factors of $\mathbb{Q}(\sqrt{-2})^{\times|C_{2}|}$ are swapped (possibly with an automorphism of $\mathbb{Q}(\sqrt{-2})$ applied depending on the choice of which factor we gave a nontrivial action to).
@@
$L\otimes_{K}L\cong\prod_{g\in G}L[x]/(x-g\alpha)\cong\prod_{g\in G}L\mathrm{.}$
@@ 5. Applications
In this section we investigate how clarification and Theorem 3.11 interact with many facets of Tambara functors. We start with a non-application.
Recall that cohomological Mackey functors may be characterized as those Mackey functors which are modules over the fixed-point Tambara functor $\mathrm{FP}(\mathbb{Z})$ corresponding to the trivial action. Cohomological Mackey functors are so-named due to the fact that group cohomology tends to provide examples of these. One might hope for some interaction between the notion of being clarified as a Green functor and being cohomological as a Mackey functor, although Example 5.1 below suggests that there isn’t anything nice to say.
@@
5.1. Affine group schemes with clarified base
In future work the author hopes to dive much deeper into the study of affine étale group schemes. For now, we give a few sample applications. Definitions and ideas in the remainder of this subsection should be viewed as protoypical rather than fixed.
Definition 5.2.
Let $k$ be a $G$-Tambara functor. An affine group scheme over $k$ is a factorization of a representable functor
$\mathrm{Hom}_{k}(R,-):k{-}\mathsf{Alg}\rightarrow\mathsf{Set}$
through the forgetful functor from the category of groups to the category of sets.
Such a factorization should endow $R$ with the structure of a cogroup object in $k$-algebras with respect to the box product over $k$. We choose this particular definition to emphasize that we are taking a functor of points perspective on Tambara functor algebraic geometry.
Proposition 5.3.
Let $k$ be a clarified $G$-Tambara functor. The representing object of any affine group scheme over $k$ has nontrivial clarification.
Heuristically, this is a statement about the failure of descent for affine group schemes along the coinduction-adjunction unit.
Proof.
The set $\mathrm{Hom}_{k}(R,k)$ is nonempty as it contains the group identity. Since $k$ is clarified, we obtain from Corollary 4.8 a morphism from the clarification of $R$ to a nonzero Tambara functor, namely, $k$. ∎
Theorem 5.4.
Let $G$ be abelian, $k$ a clarified $G$-Tambara functor with $k(G/e)$ Noetherian, and $R$ a finitely presented $k$-algebra. Then $R$ carries an action of the group $\prod_{H\subset G}G/H$ for which $G/H$ acts nontrivially if $R$ contains a factor $\mathrm{Coind}_{H}^{G}S$ in the product decomposition of Theorem 3.11. In particular, the functor
$\mathrm{Hom}_{k}(R,-)$
factors through sets with $\prod_{H\subset G}G/H$-action.
Proof.
Since $k$ is clarified and $G$ is abelian, $\mathrm{Coind}_{H}^{G}S$ carries an action of $G/H$, where $gH$ acts via the adjoint of the morphism
$\mathrm{Res}_{H}^{G}\mathrm{Coind}_{H}^{G}S\cong\prod_{gH\in G/H}S\rightarrow S$
obtained by projection onto the factor $gH$. If $S$ is nonzero, this action is nonzero, because it is given on the $G/e$ level by permuting the $G/H$ factors of $S(H/e)$ in $\mathrm{Coind}_{H}^{G}S$. These actions assemble to an action of $\prod_{H\subset G}G/H$ on any product $\prod_{H\subset G}\mathrm{Coind}_{H}^{G}S_{H}$. By Theorem 3.11, every $R$ satisfying the hypotheses of the statement of this theorem admits such a decomposition.
The last statement follows from the fact that if an object of a category admits an action of a group, then so does the $\mathrm{Hom}$-functor it represents. ∎
It is unclear to the author how Theorem 5.4 ought to interact with the notion of group scheme. Indeed, it seems unlikely that this group action on $\mathrm{Hom}$-sets lifts to a group action by group homomorphisms in the case of a group scheme, although this suggests that the definition of group scheme we propose may not be the correct one. Additionally, Lemma 2.39 suggests that coinduction of Tambara functors might be non-unitally oplax monoidal (the issue is that coinduction of Tambara functors is not left adjoint to restriction, so that lax monoidality of restriction does not imply oplax monoidality of coinduction). In other words, Proposition 5.3 seems to be the only obstruction to the problem of descent along the coinduction-adjunction unit for affine group schemes.
Switching gears, one would hope that the action of Theorem 5.4 may be made somewhat canonical; it clearly depends on how unique the isomorphism of Theorem 3.11 is (which we have some control over by Proposition 4.10). This action seems to provide a $G/H$ action on the category of algebras over $\mathrm{Coind}_{H}^{G}S$, whose fixed points form the category of $S$-algebras. In any case, it is yet more evidence that clarified Tambara functors are the most appropriate bases over which to do Tambara functor algebraic geometry, at least from the functor of points perspective.
5.2. A lift to $G$-spectra
In this subsection we discuss the topological consequences of Theorem 3.11. This discussion should be viewed more as a sketch rather than a rigorous argument, mainly owing to its reliance on results which are expected to be true but which to the author’s knowledge have not yet been proven (beyond the $G=C_{p}$ case in [Yan25]).
Let $R$ be a $G$-$\mathbb{E}_{\infty}$ ring spectrum. First defined in [LMSM86], the slightly more modern perspective is that $G$-$\mathbb{E}_{\infty}$ rings are the $\mathcal{O}$-algebras over the $N_{\infty}$ operads $\mathcal{O}$ whose corresponding indexing system is complete. Brun observed in [Bru07] that $\underline{\pi}_{0}(R)$ is a $G$-Tambara functor. Under the mild assumption that $\pi_{0}^{e}(R)$, the zeroth homotopy group of the underlying $\mathbb{E}_{\infty}$ ring of $R$, is Noetherian, Theorem 3.11 implies that $\underline{\pi}_{0}(R)$ admits a decomposition as a product of coinductions.
Suppose $\underline{\pi}_{0}(R)\cong R_{1}\times R_{2}$ for Tambara functors $R_{i}$. Let $d_{i}\in\underline{\pi}_{0}(R)(G/G)$ denote the idempotents corresponding to projection to each factor, so $d_{i}$ maps to $1$ in $R_{i}(G/G)$. Then $d_{i}$ is represented by a map $\mathbb{S}\rightarrow R$, where $\mathbb{S}$ denotes the $G$-equivariant sphere spectrum. Now we form the composite
$f_{i}:R\cong\mathbb{S}\wedge R\xrightarrow{d_{i}\wedge\mathrm{Id}_{R}}R\wedge R\rightarrow R$
where the last map is the multiplication map.
The results of [Lur17, Section 4.8.2] imply that we have an equivalence $R\cong R/f_{1}\times R/f_{2}$ of $\mathbb{E}_{\infty}$-ring spectra. Presumably the Hill-Hopkins-Ravenel norms [HHR16] respect this decomposition, so that we have a normed $\mathbb{E}_{\infty}$-algebra in the sense of Yang [Yan25]. Roughly speaking, this is an $\mathbb{E}_{\infty}$ object of $\mathrm{Sp}^{BG}$ equipped with certain maps involving Hill-Hopkins-Ravenel norms which make certain diagrams commute only up to homotopy. Yang only defines such objects for $G=C_{p}$, but it is expected that a similar definition for general finite groups $G$ should yield a similar main theorem: the category of normed $\mathbb{E}_{\infty}$-algebras in $G$-spectra is equivalent to the category of $G$-$\mathbb{E}_{\infty}$-algebras. In other words, homotopy commutativity of appropriate diagrams actually implies the a priori much stronger condition of homotopy coherence.
We emphasize that this sketch suggests that the norms of a $G$-$\mathbb{E}_{\infty}$ ring spectrum provide sufficient structure to lift product splittings of the underlying $\mathbb{E}_{\infty}$ rings in $\mathrm{Sp}^{BG}$ to product splittings of genuine $G$-spectra.
If this argument goes through, then Theorem 3.11 reduces the study of $G$-$\mathbb{E}_{\infty}$-rings $R$ with $\pi_{0}^{e}(R)$ Noetherian to the study of $G$-$\mathbb{E}_{\infty}$-rings $R$ with $\underline{\pi}_{0}(R)$ coinduced. If $\underline{\pi}_{0}(R)\cong\mathrm{Coind}_{H}^{G}k$ for some $H$-Tambara functor $k$, then $\pi_{0}(\mathrm{Res}_{e}^{G}R)\cong\mathrm{Fun}(G/H,k(H/e))$ is a product of rings permuted transitively by the $G$-action. Then we expect to again see from [Lur17, Section 4.8.2] that the underlying $\mathbb{E}_{\infty}$ object in $\mathrm{Sp}^{BG}$ of $R$ is coinduced from $H$, i.e. of the form $G/H_{+}\wedge T$ for some $\mathbb{E}_{\infty}$ object $T$ in $\mathrm{Sp}^{BH}$. As before, we may hope for this identification to lift to one of $G$-$\mathbb{E}_{\infty}$-ring spectra. In this case, the main results of [BDS15] imply that everything we might want to know about $R$ may be straightforwardly deduced from information about $T$.

### v2

@@ 4.1. Morphisms between coinduction and clarification
Every free polynomial Tambara functor on finitely many generators has level $G/e$ given by a finitely generated free polynomial $\mathbb{Z}$-algebra according to [Bru05, Theorem A]. The free polynomial $k$-algebras are obtained by base-changing the free polynomial Tambara functors along the unit map from the Burnside Tambara functor to $k$.
@@
$\prod_{H\subset G}\mathrm{Coind}_{H}^{G}R_{H}\rightarrow\prod_{H\in\Lambda}\mathrm{Coind}_{H}^{G}R_{H}.$
@@
$\prod_{g\in H\backslash G/H}\mathrm{Coind}_{H\cap{}^{g}H}^{H}\mathrm{Res}_{H\cap{}^{g}H}^{{}^{g}H}{}^{g}R\cong\mathrm{Res}_{H}^{G}\mathrm{Coind}_{H}^{G}R\rightarrow S.$
@@
We observe that each item is clarified in order. First, $G$ acts trivially on the $G/e$ level of the first three kinds Tambara functors on the list, so they are clarified. Next, The fixed-point Tambara functor associated to a Galois field extension has bottom level a field, hence the only idempotents are $1$ and 0, which are fixed by $G$. That free polynomial algebras are clarified over a clarified base is the content of Lemma 4.4. Next, we compute $\mathrm{H}^{2*}(e;\mathrm{Z})\cong\mathrm{Ext}_{\mathbb{Z}}^{2*}(\mathbb{Z},\mathbb{Z})$ which is just $\mathbb{Z}$ in degree zero, and $\underline{\pi}_{0}(MU_{G})(G/e)\cong\pi_{0}(MU)\cong\mathbb{Z}$. Any Tambara functor $k$ with $k(G/e)=\mathbb{Z}$ is clarified, so we deduce that the last two list items are clarified.
@@
Another natural and important source of coinduced Tambara functors are the free modules on a single generator (c.f. Proposition 5.2). Specifically, if $k$ is a Tambara functor, then the free module over the Green functor underlying $k$ on a single generator in level $G/H$ arises as the $k$-module structure on $\mathrm{Coind}_{H}^{G}\mathrm{Res}_{H}^{G}k$ induced by the adjunction unit
$k\rightarrow\mathrm{Coind}_{H}^{G}\mathrm{Res}_{H}^{G}k.$
In this situation, we really only needed the Green functor structure on the target, although this adjunction unit is a morphism of Tambara functors.
@@
$R\rightarrow\prod_{L\in\Lambda_{H}}\mathrm{Coind}_{L}^{G}R_{L}.$
@@
$\mathrm{Coind}_{L}^{G}(R_{L}(L/e))\rightarrow\phi_{H}(d_{L})\mathrm{Coind}_{H}^{G}R_{H}(H/e).$
@@ 4.2. Clarification as a localization
The subcategory $G{-}\mathrm{Tamb}_{\Lambda{-}\mathrm{clar}}$ of $\Lambda$-clarified $G$-Tambara functors is a reflective subcategory of the category of $G$-Tambara functors, i.e. the inclusion has a left adjoint.
@@
$R\cong\prod_{H\notin\Lambda}\mathrm{Coind}_{H}^{G}R_{H}.$
@@
Proposition 4.13 is a precise formulation of the slogan $\Lambda$-clarification is given by setting the image of coinduction equal to zero”. If $G=C_{p^{n}}$ the only clarification functors available to us are the $\Lambda_{C_{p^{m}}}$-clarifications, and a $C_{p^{n}}$-Tambara functor maps to zero under such a clarification if and only if it is in the image of coinduction from $C_{p^{m}}$-Tambara functors (assuming the $C_{p^{n}}/e$ level is Noetherian).
@@
We give an application to étale theory. We define a morphism of $G$-Tambara functors $k\rightarrow R$ to étale if $R$ is a finitely presented $k$-algebra and $k\rightarrow R$ is formally étale in the sense of [Hil17] (i.e. $R$ is flat as a $k$-module, and the genuine Kähler differentials $\Omega^{1,G}_{R/k}$ vanish).
@@
This follows from Theorem 3.10 and the fact that a product of $k$-algebras is étale if and only if each factor is. For the reader’s convenience, we sketch the argument. Write $R\cong\widehat{R}\times R^{\mathrm{clar}}$. If $R$ is finitely presented and flat over $k$, then it is straightforward to show that the same holds for $R^{\mathrm{clar}}$. Finally, one may show that the (Tambara) genuine Kähler differentials of $R$ over $k$ contain the genuine Kähler differentials of $R^{\mathrm{clar}}$ as a summand (indeed, the functor sending a $k$-algebra $R$ to the genuine Kähler differentials $\Omega^{1,G}_{R/k}$ preserves products), hence the vanishing of the former implies vanishing of the latter. ∎
@@ 4.3. Clarified colimits
$\Lambda$-clarification preserves filtered colimits. In other words, filtered colimits in $G{-}\mathrm{Tamb}_{\Lambda{-}\mathrm{clar}}$ are computed levelwise.
@@
Recall that filtered colimits of Tambara functors are computed levelwise. Since colimits in the category of $\Lambda$-clarified objects are computed by applying $\Lambda$-clarification to the corresponding colimit in $G{-}\mathrm{Tamb}$, it suffices to show that any colimit in $G{-}\mathrm{Tamb}$ of a diagram whose objects are all $\Lambda$-clarified $G$-rings is $\Lambda$-clarified.
Let $F:I\rightarrow G{-}\mathrm{Ring}$ be any filtered diagram and let $R$ be the colimit. Suppose $d\in R$ is an idempotent of type $H$ for some $H\subset G$. Then, passing to a large enough index $i\in I$, $d$ is in the image of some $\delta\in F(i)$ under canonical map $F(i)\rightarrow R$. We see that, modulo the kernel of $F(i)\rightarrow R$, $\delta$ satisfies the requisite relations to be a type $H$ idempotent.
Observe that being a type $H$ idempotent amounts to satisfying a finite list of relations: hdd=0hd-d=0 for all $h\in H$, $d\cdot gd=0$ for all $gH\in G/H$, and $d^{2}-d=0$. Thus there exists $j\in I$ and $i\rightarrow j$ such that the induced map $F(i)\rightarrow F(j)$ carries each relation $h\delta-\delta$, $\delta\cdot g\delta$, and $\delta^{2}-\delta$ to zero in $F(j)$ (because these elements map to zero in the colimit $R$). The image of $\delta$ in $F(j)$ is therefore a type $H$ idempotent.
@@
Note that $\sqrt{-2}\otimes 1+1\otimes\sqrt{-2}\mapsto(2\sqrt{-2},0)$. If we give both factors in $\mathbb{Q}(\sqrt{-2})\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-2})\cong\mathbb{Q}(\sqrt{-2})^{\times|C_{2}|}$ the trivial action, this element is fixed, and if we give both factors the Galois action, then the generator for $C_{2}$ acts by 1-1. However, if we give one factor the trivial action and the other factor the Galois action, then we observe that the factors of $\mathbb{Q}(\sqrt{-2})^{\times|C_{2}|}$ are swapped (possibly with an automorphism of $\mathbb{Q}(\sqrt{-2})$ applied depending on the choice of which factor we gave a nontrivial action to).
@@
$L\otimes_{K}L\cong\prod_{g\in G}L[x]/(x-g\alpha)\cong\prod_{g\in G}L.$
@@ 5. Applications
In this section we investigate how clarification and Theorem 3.10 interact with some facets of Tambara functors.
To start with, we note a few applications which appear in other articles. [Wis25a] studies finite affine étale group schemes; coinductions give “new” examples of finite étale extensions which are minimal in the appropriate sense. If $k$ is a Tambara functor with $k(G/e)$ an algebraically closed field and $\ell(G/e)$ is a finite étale $k$-algebra, then $k(G/e)\rightarrow\ell(G/e)$ is finite étale. Theorem 3.10 applies, showing that $\ell$ is a finite product of coinductions of Tambara functors with $G/e$-level isomorphic to the algebraically closed field $k(G/e)$.
On the other hand, [Wis25c] observes that coinduction induces an order-preserving bijection on Green ideals and Tambara ideals which restricts to a bijection of prime ideals. Consequently the adjunction units $k\rightarrow\mathrm{CoInd}_{H}^{G}\mathrm{Res}_{H}^{G}k$ lead to a stratification of the Nakaoka spectrum of $k$.
One may also show that Theorem 3.10 is suitably compatible with the transversal Tambara functors of Sulyma [Sul23]. Additionally, assuming an analogue of the main result of [Yan25] for all finite groups, one expects an analogue of Theorem 3.10 to hold for $G$-$\mathbb{E}_{\infty}$-rings $R$ under the assumption that $\pi_{0}^{G}(R)$ is a Noetherian ring.
Finally, we conclude with a non-application. Recall that cohomological Mackey functors may be characterized as those Mackey functors which are modules over the fixed-point Tambara functor $\mathrm{FP}(\mathbb{Z})$ corresponding to the trivial action. Cohomological Mackey functors are so-named due to the fact that group cohomology tends to provide examples of these. One might hope for some interaction between the notion of being clarified as a Green functor and being cohomological as a Mackey functor, although Example 5.1 below suggests that there isn’t anything nice to say.